# Decay-protected superconducting qubit with fast control enabled by integrated on-chip filters

### v1

@@ Discussion
In this work, we addressed the seemingly competing challenges of implementing brief single-qubit gates and mitigating the decoherence of qubits due to the strong coupling needed for such brief gates. These challenges are intimately related to the power dissipation budget in the dilution refrigerator that houses the qubits and cross-talk mitigation between the qubits.
A transmon qubit with a standard capacitive coupling to its broadband drive line resulting in $T_{1}$ of 1 ms requires a peak power of $-11$ dBm at room temperature to implement 10-ns-long single-qubit gates using a resonant drive. This power requirement is estimated considering 60 dB attenuation in the drive line to keep thermal qubit excitations low. In contrast, a subharmonic drive requires roughly $25$ dBm in this scenario with no drive line filter. Such drive power leads to heat dissipation of about $-17.5$ dBm at the base plate, unfortunately matching its cooling capacity of $-17$ dBm.
Our proposed on-chip filters are designed to achieve a $T_{1}$ bound of 7161 ms for the $\lambda/4$ filter and 2421 ms for the $\lambda/2$ filter at the stopband near the qubit transition frequency. We estimate peak powers of $-16$ dBm for the $\lambda/4$ filter and $-22$ dBm for the $\lambda/2$ filter at room temperature to implement 10-ns-long gates using subharmonic driving. With 60 dB of attenuation in the drive lines, the heat dissipation at the base plate is reduced to $-58$ dBm for the $\lambda/4$ filter and to $-64$ dBm for the $\lambda/2$ filter, enabling tens of thousands of simultaneous gate operations before reaching the cooling power of typical refrigerators.
@@
In our subharmonic-drive experiments, we obtained qubit spectroscopy akin to those with resonant driving using standard two-tone measurements. We measured $\pi$-pulse lengths of 37 ns and 12 ns for the $\lambda/4$ and $\lambda/2$ filters, respectively, using Rabi measurements. This result was obtained with 62 dB of attenuation distributed across various dilution stages, combined with an additional 12 dB of attenuation from filters and wires at room temperature and inside the refrigerator, yielding approximately 74 dB in total in the drive line. This configuration effectively reduced thermal noise photons to below $2.5\times 10^{-3}$ at qubit frequency and around $0.03$ at subharmonic frequency. This is a notable improvement compared to the work in Ref. [56], where single-qubit gate of length 35 ns has been demonstrated with only 43 dB attenuation in the qubit control line and without the use of bulky off-chip low-pass filters.
The off-resonant nature of the subharmonic drive led to a substantial AC Stark shift of around $-73$ MHz at a Rabi frequency of about 43 MHz with the $\lambda/2$ filter, significantly more pronounced than in resonant driving. As a result, the standard approach for implementing and characterizing single-qubit gates through Randomized Benchmarking may not be feasible as such, calling for special attention to phase correction between gates. In addition, minimization of leakage errors with such brief gates is also privotal. Although the drag scheme effectively reduces leakage errors in resonant driving, a similar approach adapted for subharmonic driving needs to consider the large AC Stark shift. However, addressing this issue falls outside the scope of this study and will be addressed in future research.
@@ Methods
where $\eta=\frac{\Omega_{\mathrm{R}}(\omega_{\mathrm{q}}-\alpha)}{\omega^{2}_{\mathrm{d}}-(\omega_{\mathrm{q}}-\alpha)^{2}}$ represents the strength of the subharmonic drive. From the Hamiltonian, we obtain the AC Stark shift $\delta_{\mathrm{sub}}=2\alpha|\eta|^{2}/3$ resulting from the off-resonant drive and the subharmonic Rabi frequency $\Omega^{\mathrm{sub}}_{\mathrm{R}}=2\alpha\eta^{3}/3$.
@@
The scattering-parameter data is exported from Sonnet to Microwave Office AWR, where we construct and simulate a full-circuit model. We extract and export the input admittance $Y_{\mathrm{in}}$ for further analysis in Python. A comprehensive description of the steps taken to achieve these simulation results is presented in the Supplementary Information S1.
@@
All reasonable data requests should be currently addressed to the corresponding author. Research data will be available on the Zenodo platform in the future.
@@
    To facilitate the design and the implementation of the on-chip filters, we establish a simulation procedure to obtain the Rabi frequency and the $T_{1}$ of the qubit resulting from the coupling to the on-chip filters. To this end, we estimate the quantum properties of superconducting circuits from classicalelectromagnetic simulations of lumped and distributed-element circuits [59, 60, 61]. By simulating the input admittance ($Y_{\mathrm{in}}$) in parallel with the qubit from qubit–environment coupling, one obtains the external coupling rate using $\gamma_{\mathrm{q}}(\omega)=\mathrm{Re}[Y_{\mathrm{in}}(\omega)]/C_{\mathrm{q}}$, where $C_{\mathrm{q}}$ is the shunt capacitance of the qubit to the ground. Here, the environment refers to any element that couples to the qubit such as standard drive lines, on-chip filters, readout resonators, and flux lines. However, care needs to be taken when addressing the contribution from the individual coupling elements, for example, by tuning out the other coupled channels by either grounding them or removing them completely from the simulation geometry.
@@
    Supplementary Fig. S1: Comparison different models for the Rabi frequency and the energy decay time of a weakly coupled qubit to a standard drive. a Simplified model of a transmon qubit coupled to a coplanar-waveguide (CPW) transmission line rendered in Sonnet. A box-wall port 1 is used at the drive port, and a co-calibrated lumped port 2 is placed on the qubit island to investigate scattering, $S_{\mathrm{21}}$, and input admittance, $Y_{\mathrm{in}}$, parameters of the qubit. Note that $Y_{\mathrm{in}}$ is calculated when the other ports are terminated using the impedances set by the port terminations, in this case, to 50 $\Omega$. An estimate of the qubit-drive coupling capacitance $C_{\mathrm{d}}$ is obtained from the finite-element method (FEM) simulation carried out in Sonnet. b Equivalent lumped model of the qubit-drive system built in Microwave Office AWR design environment. A transmission line element is used to model a CPW drive line with a characteristic impedance $Z_{\mathrm{0}}$ of 50 $\Omega$ and relative effective dielectric constant $E_{\mathrm{eff}}$ of 6.45 obtained using $E_{\mathrm{eff}}=(\epsilon_{\mathrm{Si}}+\epsilon_{\mathrm{air}})/2$, where $\epsilon_{\mathrm{Si}}=11.9$ and $\epsilon_{\mathrm{air}}=1$ are the relative dielectric permittivity of a silicon substrate and air, respectively. The coupling capacitance $C_{\mathrm{d}}$ is taken from the Sonnet simulation. c Combination of a FEM simulation and a lumped model is set up inside AWR, allowing us to simulate complex and large geometries while avoiding the time and memory constraints usually incurred in FEM simulations of such geometries. From Sonnet simulation, we obtain an S-parameter file in touchstone format and export it to AWR, where FEM-lumped-model simulation yields $Y_{\mathrm{in}}$ in parallel with the qubit due to the coupling to the drive. d, e Comparison of the estimated (d) Rabi frequency and (e) $T_{1}$ of the qubit from lumped, FEM-lumped, and analytical approaches.
@@
    Once the external coupling rate is known, the Rabi frequency and the $T_{1}$ can be estimated using $\Omega_{\mathrm{R}}(\omega)=2\sqrt{\gamma_{\mathrm{q}}(\omega)}\beta$ and $T_{1}(\omega)=1/\gamma_{\mathrm{q}}(\omega)$, respectively, as discussed in the Methods of the main text. Using this approach, we compare the Rabi frequency and the $T_{1}$ estimates obtained from FEM and lumped-element simulations to the theoretical model for a qubit weakly coupled to a standard transmission line, as shown in Supplementary Fig. S1. We observe a very good correspondence between the simulated and the theoretical results, confirming the validity of the established approach.
@@
    In Supplementary Fig. S2, we describe the approach for full-circuit simulation for the qubit Q3 coupled to the $\lambda/2$ filter, the readout resonator, and the flux line. Supplementary Figure S2a shows a simplified model of a large simulation geometry by trimming out the meanders of the $\lambda/2$ filter and adding co-calibrated lumped ports that allow for building the full-circuit scheme with a lumped-element network simulator such as Microwave office AWR. After building the full-circuit scheme in AWR, we simulate the input admittance in parallel with the qubit from the port of interest. To estimate the Rabi frequency and $T_{1}$ just from the on-chip filter, we ground all the other co-calibrated ports except the qubit and the filter ports.
@@
    Supplementary Fig. S2: Full-circuit simulation scheme. a, b (a) Simplified model of the $\lambda/2$ filter of (b) is generated in Sonnet by removing the meanders and adding two co-calibrated ports 6 and 7 at both ends. Similarly, co-calibrated ports 2, 4, and 5 are added to the qubit island, readout resonator, and flux line, respectively in the Sonnet FEM simulation. Again, box-wall port 1 is used for the qubit drive through the $\lambda/2$ filter. The FEM simulation for this geometry is carried out in the frequency range of 1–10 GHz, and the resulting 7-port S-parameter file is exported. c Full-circuit simulation scheme built in AWR utilizing the S-parameter file obtained from the FEM simulations, with the rest of the circuitry implemented with lumped elements. The meanders of the $\lambda/2$ filter corresponding to a length of about 11 mm are added as a transmission line element between ports 6 and 7 of the S-parameter network. A Purcell filter together with three readout resonators R1, R2, and R3 is constructed with just the lumped components. The parameters of readout resonators and Purcell filters such as coupling capacitances, mutual inductances, and filter and resonator lengths are adjusted to match the designed values. Readout resonator R3 is coupled to the qubit Q3 at port 4 of the S-parameter network, while the flux line at port 5 is grounded. To estimate the Rabi frequency and the $T_{1}$ from just the $\lambda/2$ filter, we also ground the resonator port, illustrated in (a) with a switch symbol.
@@
    To investigate the frequency splitting of the stopband observed in the characterization of the $\lambda/2$ filter in the experiments, we simulate a lumped model of a $\lambda/2$ filter that couples capacitively to qubit port at two separate locations, as depicted in Supplementary Fig. S3. We place a transmission line element between the coupling points, as shown in the inset of Supplementary Fig. S3b. We introduce a top and a bottom capacitor with capacitances $C^{\mathrm{top}}_{\mathrm{d}}$ and $C^{\mathrm{bot}}_{\mathrm{d}}$, respectively, at the coupling points. By adding an asymmetry offset $\Delta C^{\mathrm{top}}_{\mathrm{d}}$ to the top capacitor such that $C^{\mathrm{top}}_{\mathrm{d}}=C^{\mathrm{bot}}_{\mathrm{d}}-\Delta C^{\mathrm{top}}_{\mathrm{d}}$, we observe a frequency splitting of the stopband as a function of the offset relative to the bottom capacitor. The observed splitting of about 1 GHz in the experiment corresponds to a relative asymmetry offset of $4.9$% as shown in Supplementary Fig. S3a.
@@
    Supplementary Fig. S3: Stopband frequency splitting due to an asymmetric two-point coupling of the $\lambda/2$ filter. a Transmission amplitude from the qubit port to the drive line port as a function of frequency for the indicated asymmetry offset capacitances $\Delta C^{\mathrm{top}}_{\mathrm{d}}$ which are introduced to the top capacitor with capacitance $C^{\mathrm{top}}_{\mathrm{d}}=C^{\mathrm{bot}}_{\mathrm{d}}-\Delta C^{\mathrm{top}}_{\mathrm{d}}$. b Frequency splitting of the stopband as a function of the asymmetry offset obtained as the separation of the dips in the transmission amplitude in calcualtions identical to those exemplified in (a). For a relative asymmetry offset of approximately $4.9$% (dashed purple vertical line) with respect to the bottom capacitance $C^{\mathrm{bot}}_{\mathrm{d}}$, we obtain the frequency splitting of roughly 1 GHz, equal to that observed in the experiment. The inset shows the lumped model of the $\lambda/2$ filter used in (a).
@@
    Following the fabrication of the device, we wire-bonded the device to an SMA-8 sample holder and placed the device inside a magnetic shield, which is installed at the base plate of a Bluefors XLD500 dilution refrigerator. Throughout the experiment, the refrigerator maintained a temperature range of 18–24 mK. An overview of the experimental setup is presented in Supplementary Fig. S4.
@@
    Supplementary Fig. S6: Measurement sequence for on-chip-filter characterization for qubit Q2. a,b Normalized readout signal for the spectroscopy measurements for (a) the readout resonator and (b) the qubit Q2 as a function of the flux bias voltage and the probe frequency. c Measured resonator frequency (blue marker) and qubit frequency (turquoise marker) as a function of the bias voltage, obtained from experiments similar to those of (a) and (b), respectively. Analytical fit of the resonator frequency (red solid) and qubit frequency (purple solid) to the measured data. c,d,e For a fixed pulse amplitude, we implement (c) qubit spectroscopy by sweeping the probe frequency $f$ around qubit frequency $f_{\mathrm{q}}$, (d) Rabi oscillation by varying the qubit pulse length, and (e) $T_{1}$ measurement by changing an idle delay between the $\pi$ pulse of the qubit and the readout pulse. The measured (markers) data and corresponding fit (solid line) at two different flux bias voltages in (d)–(f) correspond to the sweet spot of the qubit at around 7.85 GHz (blue markers, purple line) and to the filter frequency around 5 GHz (turquoise markers, orange line). The fitting parameters are shown on top of the corresponding figures.
@@
    Resonator frequency
    $\mathrm{\omega}_{\mathrm{r}}/2\pi$
    6.8913 GHz
    6.9886 GHz
    6.7650 GHz
    Resonator linewidth
    $\mathrm{\kappa}/2\pi$
    2.076 MHz
    2.805 MHz
    10.756 MHz
    Qubit frequency
    $\omega_{\mathrm{q}}/2\pi$
    7.773 GHz
    7.852 GHz
    7.640 GHz
    Coupling strength
    $g/2\pi$
    72.455 MHz
    72.000 MHz
    68.750 MHz
    Anharmonicity
    $\alpha/2\pi$
    $-232.916$ MHz
    $-229.847$ MHz
    $-234.520$ MHz
    Dispersive shift
    $\chi/2\pi$
    $-2.300$ MHz
    $-2.025$ MHz
    $-1.978$ MHz (*)
    Relaxation time
    $T_{1}$
    7.446 $\mu$s
    0.481 $\mu$s
    0.06 $\mu$s
    Dephasing time
    $T^{\mathrm{*}}_{\mathrm{2}}$
    8.289 $\mu$s
    0.929 $\mu$s
    Hahn echo time
    $T^{\mathrm{E}}_{\mathrm{2}}$
    8.475 $\mu$s
    0.934 $\mu$s
@@ 0.06 $\mu$s
    We show a typical measurement sequence for the characterizion of the on-chip filters in Fig. S6 for qubit Q2 coupled to the $\lambda/4$ filter. We begin by determining the flux-bias-voltage-dependent frequencies of the resonator and of the qubit through the single-tone and two-tone spectroscopy measurements, respectively, as depicted in Fig. S6a, b. We fit the resonator and the qubit spectroscopic data to the analytical model of Ref. [62], as shown in Fig. S6c. With this fit, we can automate the search of the resonator and the qubit frequencies in large frequency ranges for a given bias voltage.
@@
    To measure the Rabi frequency and the qubit $T_{1}$, we first determine the DC voltage for the desired qubit flux bias according to Fig. S6c. Subsequently, we set the correct readout frequency according to the fit in Fig. S6c. Next, we again carry out qubit spectroscopy at very low power, fitting the data to a Lorentzian curve to accurately determine the qubit frequency.
@@
    Using the carefully extracted qubit frequency, we implemented Rabi oscillations and fit the data to a decaying cosine function to determine the $\pi$-pulse length. The Rabi oscillation measurements are conducted at fixed pulse voltages of 0.05 V and 0.025 V for the $\lambda/4$ and $\lambda/2$ filters, respectively. We then measure the $T_{1}$ time by applying a $\pi$ pulse, as obtained from the fit, and by varying the delay before the final readout pulse. We fit the measured data to a decaying exponential function to obtain the $T_{1}$ time and its uncertainty. The entire measurement sequence is automated, requiring only the bias voltage as an input parameter.

### v2

@@ Discussion
In this work, we addressed the seemingly competing challenges of implementing fast single-qubit gates and mitigating the decoherence of qubits due to the strong coupling needed for such fast gates. These challenges are intimately related to the power dissipation budget in the dilution refrigerator that houses the qubits and cross-talk mitigation between the qubits.
A transmon qubit with a standard capacitive coupling to its broadband drive line resulting in $T_{\mathrm{1,ext}}$ of 1 ms requires a peak power of $-11$ dBm at room temperature to implement 10-ns-long single-qubit gates using a resonant drive. This power requirement is estimated considering 60 dB attenuation in the drive line to keep thermal qubit excitations low. In contrast, a subharmonic drive requires roughly $25$ dBm in this scenario with no drive line filter. Such drive power leads to heat dissipation of about $-17.5$ dBm at the base plate, unfortunately matching its cooling capacity of $-17$ dBm.
Our proposed on-chip filters are designed to achieve a $T_{\mathrm{1,ext}}$ bound of 7161 ms for the $\lambda/4$ filter and 2421 ms for the $\lambda/2$ filter at the stopband near the qubit transition frequency. We estimate peak powers of $-16$ dBm for the $\lambda/4$ filter and $-22$ dBm for the $\lambda/2$ filter at room temperature to implement 10-ns-long gates using subharmonic driving. With 60 dB of attenuation in the drive lines, the heat dissipation at the base plate is reduced to $-58$ dBm for the $\lambda/4$ filter and to $-64$ dBm for the $\lambda/2$ filter, enabling tens of thousands of simultaneous gate operations before reaching the cooling power of typical refrigerators.
@@
In our subharmonic-drive experiments, we obtained qubit spectroscopy akin to those with resonant driving using standard two-tone measurements. We measured $\pi$ pulse lengths of 37 ns and 12 ns for the $\lambda/4$ and $\lambda/2$ filters, respectively, using Rabi measurements. This result was obtained with 62 dB of attenuation distributed across various dilution stages, combined with an additional 12 dB of attenuation from filters and wires at room temperature and inside the refrigerator, yielding approximately 74 dB in total in the drive line. This configuration effectively reduced thermal noise photons to below $2.5\times 10^{-3}$ at qubit frequency and around $0.03$ at subharmonic frequency. This is a prominent improvement compared to the work in Ref. [60], where single-qubit gate of length 35 ns has been demonstrated with only 43 dB attenuation in the qubit control line and with the use of bulky off-chip low-pass filters.
The off-resonant nature of the subharmonic drive led to a substantial AC Stark shift of around $-73$ MHz at a Rabi frequency of about 43 MHz with the $\lambda/2$ filter, significantly more pronounced than in resonant driving. As a result, the standard approach for implementing and characterizing single-qubit gates through Randomized Benchmarking may not be feasible as such, calling for special attention to phase correction between gates. In addition, minimization of leakage errors with such brief gates is also pivotal. Although the drag scheme effectively reduces leakage errors in resonant driving, a similar approach adapted for subharmonic driving needs to consider the large AC Stark shift. Addressing this issue stands as an interesting topic for future research.
@@ Methods
where $\eta=\frac{\Omega_{\mathrm{R}}(\omega_{\mathrm{q}}-\alpha)}{\omega^{2}_{\mathrm{d}}-(\omega_{\mathrm{q}}-\alpha)^{2}}$ represents the strength of the subharmonic drive. From the Hamiltonian, we obtain the AC Stark shift $\delta_{\mathrm{sub}}=2\alpha|\eta|^{2}/3$ resulting from the off-resonant drive and the subharmonic Rabi frequency $\Omega^{\mathrm{sub}}_{\mathrm{R}}=2\alpha|\eta|^{3}/3$.
@@
The scattering-parameter data is exported from Sonnet to Microwave Office AWR, where we construct and simulate a full-circuit model. We extract and export the input admittance $Y_{\mathrm{in}}$ for further analysis in Python. A comprehensive description of the steps taken to achieve these simulation results is presented in the Supplementary Note S1 in Supplementary Information.
@@
All relevant data and codes generating the figures in this article are available via Zenodo at https://doi.org/doi/10.5281/zenodo.11234842 [68].
Acknowledgments
We express our gratitude to the KAUTE Foundation for their support through the Ph.D. grant awarded to the first author. We acknowledge the European Research Council under the Advanced Grant no. 101053801 (ConceptQ), Business Finland under the Quantum Technologies Industrial project (Grant no. 2118781) and Academy of Finland under its Centre of Excellence Quantum Technology Finland (Grant no. 352925) and through the Finnish Quantum Flagship. Special thanks are extended to Sergei Lemziakov, Dmitrii Lvov, and Joonas Peltonen for their invaluable assistance with the Otanano cryogenic facility. The authors acknowledge Giacomo Catto, Timm Mörstedt, and Priyank Singh for useful discussions related to the fabrication and Jukka-Pekka Kaikkonen from VTT Technical Research for providing us with the unpatterned Niobium-on-Silicon wafer. Furthermore, we are thankful for the enriching general discussions on the subject matter with Michael Hatridge, Florian Blanchet, Arman Alizadeh, Wallace Santos Teixeira, Vasilii Vadimov, and Johannes Heinsoo. We thank Eric Hyyppä for encouraging discussions and insightful comments on the manuscript.
Authors contributions
A.S. and M.M. conceived the experiment. A.S. and S.K. designed and carried out the electromagnetic simulations of the device. H.S. and A.S. wrote the measurement codes for the experiment. Q.C. helped with the theoretical discussions on subharmonic drive. A.S. fabricated the device, conducted the experiments, and analyzed the results with feedback from S.K. A.S., S.K., and M.M. wrote the manuscript with comments from all the authors.
Competing interests
M.M. declares that he is a Co-Founder and Shareholder of IQM Finland Oy. All other authors declare no competing interests.
@@ Acknowledgments
To facilitate the design and the implementation of the on-chip filters, we establish a simulation procedure to obtain the Rabi frequency and the $T_{\mathrm{1,ext}}$ of the qubit resulting from the coupling to the on-chip filters. To this end, we estimate the quantum properties of superconducting circuits from classical electromagnetic simulations of lumped and distributed-element circuits [65, 66, 67], see Supplementary Fig. S1a-c. By simulating the input admittance ($Y_{\mathrm{in}}$) in parallel with the qubit from qubit–environment coupling, one obtains the external coupling rate using $\gamma_{\mathrm{q}}(\omega)=\mathrm{Re}[Y_{\mathrm{in}}(\omega)]/C_{\mathrm{q}}$, where $C_{\mathrm{q}}$ is the shunt capacitance of the qubit to the ground. Here, the environment refers to any element that couples to the qubit such as standard drive lines, on-chip filters, readout resonators, and flux lines. However, care needs to be taken when addressing the contribution from the individual coupling elements, for example, by tuning out the other coupled channels by either grounding them or removing them completely from the simulation geometry.
@@
Supplementary Fig. S1: Comparison different models for the Rabi frequency and the energy decay time of a weakly coupled qubit to a standard drive. a Simplified model of a transmon qubit coupled to a coplanar-waveguide (CPW) transmission line rendered in Sonnet. A box-wall port 1 is used at the drive port, and a co-calibrated lumped port 2 is placed on the qubit island to investigate scattering, $S_{\mathrm{21}}$, and input admittance, $Y_{\mathrm{in}}$, parameters of the qubit. Note that $Y_{\mathrm{in}}$ is calculated when the other ports are terminated using the impedances set by the port terminations, in this case, to 50 $\Omega$. An estimate of the qubit-drive coupling capacitance $C_{\mathrm{d}}$ is obtained from the finite-element method (FEM) simulation carried out in Sonnet. b Equivalent lumped model of the qubit-drive system built in Microwave Office AWR design environment. A transmission line element is used to model a CPW drive line with a characteristic impedance $Z_{\mathrm{0}}$ of 50 $\Omega$ and relative effective dielectric constant $E_{\mathrm{eff}}$ of 6.45 obtained using $E_{\mathrm{eff}}=(\epsilon_{\mathrm{Si}}+\epsilon_{\mathrm{air}})/2$, where $\epsilon_{\mathrm{Si}}=11.9$ and $\epsilon_{\mathrm{air}}=1$ are the relative dielectric permittivity of a silicon substrate and air, respectively. The coupling capacitance $C_{\mathrm{d}}$ is taken from the Sonnet simulation. c Combination of a FEM simulation and a lumped model is set up inside AWR, allowing us to simulate complex and large geometries while avoiding the time and memory constraints usually incurred in FEM simulations of such geometries. From Sonnet simulation, we obtain a S-parameter file in touchstone format and export it to AWR, where FEM-lumped-model simulation yields $Y_{\mathrm{in}}$ in parallel with the qubit due to the coupling to the drive. d, e Comparison of the estimated (d) Rabi frequency and (e) Relaxation time, $T_{\mathrm{1,ext}}$, of the qubit from lumped, FEM-lumped, and analytical approaches.
@@
Once the external coupling rate is known, the Rabi frequency and the $T_{\mathrm{1,ext}}$ can be estimated using $\Omega_{\mathrm{R}}(\omega)=2\sqrt{\gamma_{\mathrm{q}}(\omega)}\beta$ and $T_{\mathrm{1,ext}}(\omega)=1/\gamma_{\mathrm{q}}(\omega)$, respectively, as discussed in the Methods of the main text. Using this approach, we compare the Rabi frequency and the $T_{\mathrm{1,ext}}$ estimates obtained from FEM and lumped-element simulations to the theoretical model for a qubit weakly coupled to a standard transmission line, as shown in Supplementary Fig. S1d,e. We observe a very good correspondence between the simulated and the theoretical results, confirming the validity of the established approach.
@@
In Supplementary Fig. S2, we describe the approach for full-circuit simulation for the qubit Q3 coupled to the $\lambda/2$ filter, the readout resonator, and the flux line. Supplementary Figure S2a shows a simplified model of a large simulation geometry, depicted in Supplementary Fig. S2b, by trimming out the meanders of the $\lambda/2$ filter and adding co-calibrated lumped ports that allow for building the full-circuit scheme with a lumped-element network simulator such as Microwave office AWR. After building the full-circuit scheme in AWR, we simulate the input admittance in parallel with the qubit from the port of interest, as shown in Supplementary Fig. S2c. To estimate the Rabi frequency and $T_{\mathrm{1,ext}}$ just from the on-chip filter, we ground all the other co-calibrated ports except the qubit and the filter ports.
@@
Supplementary Fig. S2: Full-circuit simulation scheme. a, b (a) Simplified model of the $\lambda/2$ filter of (b) is generated in Sonnet by removing the meanders and adding two co-calibrated ports 6 and 7 at both ends. Similarly, co-calibrated ports 2, 4, and 5 are added to the qubit island, readout resonator, and flux line, respectively in the Sonnet FEM simulation. Again, box-wall port 1 is used for the qubit drive through the $\lambda/2$ filter. The FEM simulation for this geometry is carried out in the frequency range of 1–10 GHz, and the resulting 7-port S-parameter file is exported. c Full-circuit simulation scheme built in AWR utilizing the S-parameter file obtained from the FEM simulations, with the rest of the circuitry implemented with lumped elements. The meanders of the $\lambda/2$ filter corresponding to a length of about 11 mm are added as a transmission line element between ports 6 and 7 of the S-parameter network. A Purcell filter together with three readout resonators R1, R2, and R3 is constructed with just the lumped components. The parameters of readout resonators and Purcell filters such as coupling capacitances, mutual inductances, and filter and resonator lengths are adjusted to match the designed values. Readout resonator R3 is coupled to the qubit Q3 at port 4 of the S-parameter network, while the flux line at port 5 is grounded. To estimate the Rabi frequency and the $T_{\mathrm{1,ext}}$ from just the $\lambda/2$ filter, we also ground the resonator port, illustrated in (a) with a switch symbol.
@@
To investigate the frequency splitting of the stopband observed in the characterization of the $\lambda/2$ filter in the experiments, we simulate a lumped model of a $\lambda/2$ filter that couples capacitively to qubit port at two separate locations, as depicted in Supplementary Fig. S3. We place a transmission line element between the coupling points, as shown in Supplementary Fig. S3a. We introduce a top and a bottom capacitor with capacitances $C^{\mathrm{top}}_{\mathrm{d}}$ and $C^{\mathrm{bot}}_{\mathrm{d}}$, respectively, at the coupling points. By adding an asymmetry offset $\Delta C^{\mathrm{top}}_{\mathrm{d}}$ to the top capacitor such that $C^{\mathrm{top}}_{\mathrm{d}}=C^{\mathrm{bot}}_{\mathrm{d}}-\Delta C^{\mathrm{top}}_{\mathrm{d}}$, we observe a frequency splitting of the stopband as a function of the offset relative to the bottom capacitor, as depicted in Supplementary Fig. S3b, c. The observed splitting of about 1 GHz in the experiment corresponds to a relative asymmetry offset of $4.9$% as shown in Supplementary Fig. S3b, d. Furthermore, we conclude that the splitting of the stopband retains the simulated $T_{\mathrm{1,ext}}$ in the range of a few seconds; however, it does so with reduced bandwidth. This reduction in bandwidth occurs because the $\lambda/2$ filter no longer behaves as a high-order filter but as a first-order filter due to asymmetric coupling.
Additionally, we report the filter’s sensitivity to the asymmetric offset of the top capacitor, which results in stopband splitting within the filter’s bandwidth of approximately 500 MHz (estimated at $T_{\mathrm{1,ext}}$ of 1 ms for an ideal symmetric coupling case). Our simulations indicate that a deviation of $\Delta C^{\mathrm{top}}_{\mathrm{d}}/C^{\mathrm{bot}}_{\mathrm{d}}$ within 1.2% meets the filter sensitivity conditions.
@@
Supplementary Fig. S3: Stopband frequency splitting due to an asymmetric two-point coupling of the $\lambda/2$ filter. a Lumped model of the $\lambda/2$ filter used for the simulation of asymmetric two-point coupling. An offset capacitance $\Delta C^{\mathrm{top}}_{\mathrm{d}}$ is introduced to the top capacitor with capacitance $C^{\mathrm{top}}_{\mathrm{d}}=C^{\mathrm{bot}}_{\mathrm{d}}-\Delta C^{\mathrm{top}}_{\mathrm{d}}$. b Transmission amplitude from the qubit port to the drive line port as a function of frequency for the indicated asymmetry offset capacitances $\Delta C^{\mathrm{top}}_{\mathrm{d}}$ relative to the bottom capacitor $C^{\mathrm{bot}}_{\mathrm{d}}$. c Simulated $T_{\mathrm{1,ext}}$ due to the external coupling as a function of frequency for similar asymmetric offset as in (b). d Frequency splitting of the stopband as a function of the asymmetry offset obtained as the separation of the dips in the transmission amplitude in calculations identical to those exemplified in (b). For a relative asymmetry offset of approximately $4.9$% (dashed purple vertical line) with respect to the bottom capacitance $C^{\mathrm{bot}}_{\mathrm{d}}$, we obtain the frequency splitting of roughly 1 GHz, equal to that observed in the experiment.
@@
Following the fabrication of the device, we wire-bonded the device to a SMA-8 sample holder and placed the device inside a magnetic shield, which is installed at the base plate of a Bluefors XLD500 dilution refrigerator. Throughout the experiment, the refrigerator maintained a temperature range of 18–24 mK. An overview of the experimental setup is presented in Supplementary Fig. S4.
@@
Resonator frequency $\mathrm{\omega}_{\mathrm{r}}/2\pi$ 6.8913 GHz 6.9886 GHz 6.7650 GHz
Resonator linewidth $\mathrm{\kappa}/2\pi$ 2.076 MHz 2.805 MHz 10.756 MHz
Qubit frequency $\omega_{\mathrm{q}}/2\pi$ 7.773 GHz 7.852 GHz 7.640 GHz
Coupling strength $g/2\pi$ 72.455 MHz 72.000 MHz 68.750 MHz
Anharmonicity $\alpha/2\pi$ $-232.916$ MHz $-229.847$ MHz $-234.520$ MHz
Dispersive shift $\chi/2\pi$ $-2.300$ MHz $-2.025$ MHz $-1.978$ MHz (*)
Relaxation time $T_{1}$ 7.446 $\mu$s 0.481 $\mu$s 0.06 $\mu$s
Dephasing time $T^{\mathrm{*}}_{\mathrm{2}}$ 8.289 $\mu$s 0.929 $\mu$s
Hahn echo time $T^{\mathrm{E}}_{\mathrm{2}}$ 8.475 $\mu$s 0.934 $\mu$s
Supplementary Fig. S6: Measurement sequence for on-chip-filter characterization for qubit Q2. a,b Normalized readout signal for the spectroscopy measurements for (a) the readout resonator and (b) the qubit Q2 as a function of the flux bias voltage and the probe frequency. c Measured resonator frequency (blue marker) and qubit frequency (turquoise marker) as a function of the bias voltage, obtained from experiments similar to those of (a) and (b), respectively. Analytical fit of the resonator frequency (red solid) and qubit frequency (purple solid) to the measured data. d,e,f For a fixed pulse amplitude, we implement (d) qubit spectroscopy by sweeping the probe frequency $f$ around qubit frequency $f_{\mathrm{q}}$, (e) Rabi oscillation by varying the qubit pulse length, and (f) $T_{1}$ measurement by changing an idle delay between the $\pi$ pulse of the qubit and the readout pulse. The measured (markers) data and corresponding fit (solid line) at two different flux bias voltages in (d)–(f) correspond to the sweet spot of the qubit at around 7.85 GHz (blue markers, purple line) and to the filter frequency around 5 GHz (turquoise markers, orange line). The fitting parameters are shown on top of the corresponding figures.
@@
We show a typical measurement sequence for the characterizion of the on-chip filters in Supplementary Fig. S6 for qubit Q2 coupled to the $\lambda/4$ filter. We begin by determining the flux-bias-voltage-dependent frequencies of the resonator and of the qubit through the single-tone and two-tone spectroscopy measurements, respectively, as depicted in Supplementary Fig. S6a, b. We fit the resonator and the qubit spectroscopic data to the analytical model of Ref. [69], as shown in Supplementary Fig. S6c. With this fit, we can automate the search of the resonator and the qubit frequencies in large frequency ranges for a given bias voltage.
@@
In Supplementary Fig. S6d-e, we show a measurement sequence for two different qubit flux bias voltages, corresponding to the qubit frequencies of 7.851 GHz at the sweet-spot and 5.023 GHz at the $\lambda/4$ filter stop band. To measure the Rabi frequency and the qubit $T_{1}$, we first determine the DC voltage for the desired qubit flux bias according to Supplementary Fig. S6c. Subsequently, we set the correct readout frequency according to the fit in Supplementary Fig. S6c. Next, we again carry out qubit spectroscopy at very low power, fitting the data to a Lorentzian curve to accurately determine the qubit frequency, see Supplementary Fig. S6d.
@@
Using the carefully extracted qubit frequency, we implemented Rabi oscillations and fit the data to a decaying cosine function to determine the $\pi$ pulse length. The Rabi oscillation measurements are conducted at fixed pulse voltage of 0.05 V, as shown in Supplementary Fig. S6e. We then measure the $T_{1}$ time by applying a $\pi$ pulse, as obtained from the Rabi oscillation fit, and by varying the delay before the final readout pulse. We fit the measured data to a decaying exponential function to obtain the $T_{1}$ time and its uncertainty, see Supplementary Fig. S6f. The entire measurement sequence is automated, requiring only the bias voltage as an input parameter.